Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-page-20131110

Observers: Susanna Spiro        TO: Monica Castillo
Support team: S.Benetti, N.Elias-Rosa
TAT on duty: Joe Anderson, Matt Nicholl
------------------------------------------------

Subrun 3, night 3

Instruments: EFOSC

afternoon CALIBRATIONS
PESSTO_bias         avg~198 constant, std~10
PESSTO_calib        only spec_cal_arcs (due to bad weather it is not possible for now to move the telescope to take FF)
     

Dome closed the whole night due to ice (consequence of the bad weather conditions during the day: humidity 90-100%, temperature ~0, no wind).
Comments